OpenStudy (anonymous):

12x^2y-5xy^2-3y^3 factor, show steps please

OpenStudy (anonymous):

ignore the y for a second. you have then 12x^2 - 5x - 3 = ( 4x -3) ( 3x + 1)

OpenStudy (anonymous):

12x^2y-5xy^2-3y^3 = y ( 12x^2 - 5x y - 3y^2)

OpenStudy (anonymous):

12x^2y-5xy^2-3y^3 = y ( 12x^2 - 5x y - 3y^2) = y ( 4x -3y) ( 3x + y)